OpenStudy (anonymous):

Solve the system of equations and choose the correct ordered pair x + 4y = -10 4x - 3y = -2

OpenStudy (anonymous):

looking for x and y?

OpenStudy (anonymous):

yea so what would it be

OpenStudy (anonymous):

x=-10-4y sub x into 4x-3y=-2 4(-10-4y)-3y=-2 -40-16y-3y=-2 -19y=38 y=-2 sub y=-2 into x+4y=-10 x+4(-2)=-10 x=-2

OpenStudy (anonymous):

whats the ordered pairs

OpenStudy (anonymous):

what do you mean by the ordered pairs?

OpenStudy (anonymous):

I believe what it means is that there are two answers....one for x and one for y....i.e. a pair. Then the pair would be x=-2 and y = -2.